A fan motor uses 8310 J of energy in 9.44 s. This energy is used to accelerate a 8.12 kg blade to 19.7m/s in 3.10 s. What is the efficiency of the fan? Leave your answer as a decimal.


1
Expert's answer
2021-06-29T08:17:45-0400

Solution.

"W_m=8310J;"

"t_m=9.44s;"

"m=8.12kg;"

"v=19.7m\/s;"

"t=3.10s;"

"W_K=\dfrac{mv^2}{2};"

"W_K=\dfrac{8.12\sdot19.7^2}{2}=1575.6J;"

"P_{out}=\dfrac{W_K}{t};"

"P_{out}=\dfrac{1575.6}{3.10}=508.3w;"

"P_{in}=\dfrac{8310}{9.44}=880.3w;"

"efficiency=\dfrac{P_{out}}{P_{in}};"


"efficiency=\dfrac{508.3}{880.3}=0.58;"


Answer:"efficiency=0.58;"